OpenStudy (anonymous):

Which of the following are a possible solution to the (3x - 1)(5x + 20) = 0? A. 1/3, 4 B. 1/3, -4 C. 1, 20 D. -1, -20

OpenStudy (anonymous):

B

Directrix (directrix):

(3x - 1)(5x + 20) = 0 By the Zero Product Property, 3x - 1 = 0 or 5x + 20 = 0 3x = 1 or 5x = - 20 x = 1/3 or x = -4 ---> B